Assume that the aggregate production of an economy is Yt= √Kt Lt , where Kt+1 =( 1-d) Kt+ It, St= sYt and Lt= L( i.e., the notation and meanings is to correspond the setting for the allow model with constant population. Then the savings rate s that maximizes the steady rate of consumption equals ?
1
Expert's answer
2020-05-05T18:20:06-0400

"ct=(1-s)\times Y t"

"s=\frac{St}{Yt}"

"Yt=\sqrt{Kt}\times Lt"

"Kt + 1 = (1-d)\times Kt + It"

"Kt = (1-d)\times Kt+It-1"

Lt = L

"Yt=\sqrt{((1-d) Kt + It-1)}\times L"

"ct=(1-\frac{St}{Yt})\times Yt=Yt-St=\sqrt{(1-d) Kt + It-1)}\times L-St"